Ex 3.2 Class 6 Maths Question 2.
State whether the following statements are True or False.
(a) The sum of three odd numbers is even.
(b) The sum of two odd numbers and one even number is even.
(c) The product of three odd numbers is odd.
(d) If an even number is divided by 2, the quotient is always odd.
(e) All prime numbers are odd.
(f) Prime numbers do not have any factors.
(g) Sum of two prime numbers is always even.
(h) 2 is only the even prime number.
(i) All even numbers are composite numbers. (j) The product of any two even numbers is always even.

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(а) False [∵ 3 + 5 + 7 = 15 (odd)] (b) True [∵ 3 + 5 + 6 = 14 (even)] (c) True [∵ 5 x 7 x 9 = 315 (odd)] (d) False [∵ 6 + 2 = 3 (odd)] (e) False [∵ 2 is a prime number but it is even] (f) False [∵ 3 is a prime number having 1 and 3 as its factors] (g) False [∵ 7 + 2 = 9 (odd)] (h) True [∵ 2 is even and the lowest prime number] (i) False [∵ 2 is even but not composite number] (j) True [∵ 4 x 6 = 24 (even)]